The following annual per unit cost data at an activity level of 10,000 units has been provided below. the company produces and sells only one product.

($ per unit)
sales 150.00
direct labor 15.00
direct material 18.00
variable manufacturing overhead 7.50
fixed manufacturing overhead 50.00
variable selling expense 1.50
fixed selling expenses 40.00
fixed administrative expenses 15.00

the relevant range is 8,000-13,000
at the current level of activity, calculate the following:

total sales revenue:
total contribution margin:
total fixed expenses:

the company is considering a price reduction approach to increase net income. the company believes that a 7% reduction in sales price will increase volume by 15%.
simultaneously there will be a change in suppliers for direct material that will result in lowering direct materials costs by 10%.
calculate the following:
total sales revenue:
total contribution margin:
total fixed expenses:
Answer #1

Sales revenue = Number of units sold x Selling price per unit

= 10,000 x 150

= $1,500,000

Variable cost per unit = Direct materials + Direct labor + Variable manufacturing overhead + Variable selling expenses

= 18+15+7.5+1.5

= $42

Total variable cost = Variable cost per unit x Number of units sold

= 42 x 10,000

= $420,000

Total contribution margin = Sales revenue - Total variable cost

= 1,500,000-420,000

= $1,080,000

Total fixed expenses = Fixed manufacturing overhead + Fixed selling expenses + Fixed administrative expenses

= 10,000 x 50 + 10,000 x 40 + 10,000 x 15

= 500,000+400,000+150,000

= $1,050,000

Kindly comment if you need further assistance. Thanks‼!